Define h(0,0) in such a way that h is continuous at origin h(u,v)=ln⁡((nu^2-u^2 v^2+nv^2)/(u^2+v^2 ))



1
Expert's answer
2021-05-07T09:37:20-0400

"h(u,v)=\\ln\u2061\\dfrac{nu^2-u^2 v^2+nv^2}{u^2+v^2 }"


"h(u,v)=\\ln\u2061\\dfrac{nu^2+nv^2-u^2 v^2}{u^2+v^2 }"


"h(u,v)=\\ln\u2061\\dfrac{n(u^2+v^2)-(uv)^2}{u^2+v^2 }"


"h(u,v)=\\ln\u2061(\\dfrac{n(u^2+v^2)}{u^2+v^2 }-\\dfrac{(uv)^2}{u^2+v^2 })"


"h(u,v)=\\ln(\u2061{n}- \\dfrac{(uv)^2}{u^2+v^2 })"


"\lim\limits_{u,v \to 0,0}= \ln n"


"\therefore" h(u,v) is not continuous at origin.